# Spatially resolved H emission in B14-65666: compact starbursts, ionizing efficiency and gas kinematics in an advanced merger at the Epoch of Reionization

### v1

@@ 4.5 Kinematics of the ionized ISM
The kinematics of the ionized gas can be measured with the H emission line profile in our MRS data. Based on the integrated H line, we measure a redshift of â=â7.1513ââ0.0007 for the B14-65666 system, compatible with the values obtained with the FIR emission lines [OâIII]â88âm and [CâII]â158âm (â=â7.1521ââ0.0004, Hashimoto et al. 2019), and optical [OâIII] line (â=â7.1517ââ0.0001, Jones et al. 2024). The two galaxies of the B14-65666 system have a relative velocity of 175ââ28âkmâs-1 along our line of sight as derived from the centroid of the H line emission, showing it is a galaxy merger.
@@
The kinematics of the ionized gas traced by H appears to show differences between the two galaxies. Specifically, we find that the FWHM for galaxy W is higher than for galaxy E, with values of 312ââ44âkmâs-1, and 243ââ41âkmâs-1, respectively. These values agree within the uncertainties from those measured with the far-infrared emission lines, [CâII]â158âm and [OâIII]â88âm. The [CâII]â158âm emission has similar kinematics in both galaxies (FWHM of 288ââ41âkmâs-1 for galaxy W and 276ââ34âkmâs-1 555The FWHM value for clump E has been recalculated and differs from that reported in Hashimoto et al. (2019), as confirmed through private communication with the authors. for galaxy E). For [OâIII]â88âm, galaxy E shows a broader line (322ââ41âkmâs-1) than galaxy W (267ââ42âkmâs-1). In Fig. 6 we present the comparison of these three lines for galaxies E+W and E and W. The observed differences in the two galaxies indicate the complex structure of the velocity field in the different phases of the interstellar medium. While the H line traces the overall kinematics of the ionized gas, the [CâII]â158âm line is closely associated with the neutral gas and the [OâIII]â88 m with the high excitation gas. Simulations also show that different tracers are sensitive to different kinematic properties (Kohandel et al. 2024). In addition, the continuum imaging (Sugahara et al. 2025) of both galaxies shows a different morphology, galaxy E as a very compact source while galaxy W shows a more extended structure, in particular in the rest-ultraviolet light, reminiscent of tidal tails. Moreover, low spectral resolution (Rââ1000) integral field spectroscopy with NIRSpec suggests the presence of a narrow and a broad kinematic component in both galaxies and in all optical emission lines tracing the ionized gas (Jones et al. 2024). These two components are potentially associated with the motions in the host galaxy (narrow), and the presence of high-velocity gas outflows (broad). The width of the narrow component is only 144ââ11âkmâs-1 for galaxy E and 227ââ6âkmâs-1 for galaxy W. These are significantly narrower than the values derived from our H line and far-infrared emission lines, obtained with a better spectral resolution but lower SNR. We test whether the observed H line profile could be compatible with the combination of the narrow and broad (653ââ10âkmâs-1 for galaxy E and 755ââ23âkmâs-1 for galaxy W) line components identified with NIRSpec. Our tests (see Fig. 7) indicate that the narrow+broad (N+B) component could be slightly compatible with our observed H line profiles. However, the simulated H N+B line profile appears to have a flux excess in the spectral range of the line (i.e. residuals mostly below zero) for galaxy E while the faint broad component in galaxy W is close in velocity to the peak emission in galaxy E. Since the two galaxies are separated by only 0.44 arcsec, this secondary emission in W could be due to a contamination from the brighter emission from E (Jones et al. 2024). Further higher spectral resolution data with NIRSpec, and higher SNR MRS spectra, would be required to confirm the presence of such a broad line emission in the two galaxies.
@@
The consistency of the trends between our results and the ancillary data provides robust evidence of the dynamic differences between galaxies E and W. The broader profiles observed in galaxy W for both H and [CâII]â158âm may indicate enhanced turbulence or kinematic effects, potentially driven by interactions or feedback processes, as well as a possibility of a connection with outflows. On the other hand, the higher FWHM in [OâIII]â88 m for galaxy E suggests that highly ionized gas dynamics differ from those of lower-ionization states, pointing to localized physical conditions or energetic phenomena. Also, dust attenuation could impact the interpretation, in particular in galaxy E where a large visual extinction is measured. The optical lines are more susceptible to dust extinction, which can affect measurements of line widths and kinematic features, whereas far-infrared lines are less impacted by such effects.
@@ 4.6 Stellar and dynamical mass
Although the line profiles of the different emission lines suggest that the gas is not fully (dynamically) relaxed, here we obtain an upper limit of the dynamical mass of each of the galaxies assuming the system is close to relaxation and characterized by the random motions traced by the ionized gas velocity dispersion, i.e. not dominated by rotational motions. Under this working hypothesis, we follow previous work in similar systems like the low-z luminous infrared galaxies (e.g. equation (1) in Bellocchi et al. 2013). The dynamical masses we derive are (1.5ââ0.3)âââ and (8.5ââ2.4)âââ for galaxies E and W, respectively assuming velocity dispersions of 103ââ17 kmâs-1 and 132ââ19âkmâs-1 (obtained with the FWHM from our H line), and the effective radii derived from our NIRCam images (see Sect. 3.4 for details). The total stellar masses are derived from the multi-wavelength SED fitting and correspond to (15ââ9)âââMâ and (8ââ4)ââ108âMâ for galaxies E and W. Thus, while there is a good agreement between the stellar and dynamical mass for galaxy E, an order of magnitude difference appears for galaxy W. The difference in galaxy W could be due to the presence of a dominant molecular gas component in this galaxy amounting to close to 90% of the total mass. However, the stellar structure shows the presence of extended emission consisting of star-forming clumps along tidal tails (Sugahara et al. 2025). This suggests that this galaxy is far from relaxed and therefore the estimated dynamical mass is highly uncertain.
@@ 4.7 Mass-size relation
The two galaxies of the B14-65666 system occupy a very different location in the stellar mass-size plane (see Fig. 8). Galaxy W with an effective radius of 348 pc and a stellar mass of (8ââ4) 108 Mâ, is consistent with the expected values according to the relations derived for high redshift galaxies from the JADES program (redshifts 4 to 10; Langeroodi & Hjorth 2023), and from the CEERS, PRIMER-UDS, and PRIMER-COSMOS surveys (redshifts 3 to 9; Allen et al. 2025). Therefore, galaxy W appears as an average galaxy at redshift of about 7. On the other hand, the second component of the system, galaxy E, appears as deviating from the standard relations, showing a very small, compact size (upper limit of 63 pc to the effective radius) for a stellar mass of 15 (â9)ââ108âMâ. This translates into a high stellar mass surface density (6âââMââpc-2 for E, versus 1â Mââpc-2 for W), close to the limit of the highest values (âMââpc-2) measured in the stellar clusters and nuclei of low-z galaxies (GrudiÄ et al. 2019). Galaxies as compact as galaxy E and with similar stellar surface densities are not uncommon at high- (Ălvarez-MĂĄrquez et al. 2025). Star-forming clumps and galaxies with effective radius in the 30 to 200 pc range and a wide range of mass (4ââ107 to 2ââ1010 Mâ) are already known at redshifts above 6 (see Fig. 8; Claeyssens et al. 2023; Baggen et al. 2023).
Thus, B14-65666 is identified as a system that consists of a pair of interacting galaxies, with similar stellar masses (within a factor of 2) but with one of the galaxies (E) being extremely compact and already showing stellar mass surface densities close to those measured in the nuclei of galaxies at low redshift. Even more, the galaxies of the B14-65666 system are much less massive (ââ20âââ30), and far more compact (ââ10âââ60) than the main-sequence and the most luminous star-forming galaxies at low redshift, with effective radius â3.5 kpc (Leroy et al. 2021), and â2.2 kpc (Bellocchi et al. 2022) respectively (see Fig. 8). Thus, the galaxies E and W appear as building up their stellar mass as they go through the advanced phases of the interaction/merging process.
@@ 4.8 Starburst nature
The analysis of the SFR derived with H versus stellar mass plane reveals that both galaxies E and W are well above the main sequence (MS) of galaxies at redshift 7 (Heintz et al. 2023). Galaxies E and W of our system are situated between the region defined by Caputi et al. (2017) for H emitters, the starburst region (SB), and the MS (see Fig. 9). Such extremely compact starburst galaxies are expected to have short depletion times (¥100 Myr, see Franco et al. 2020) and therefore galaxy E may quench soon out of gas exhaustion - at least temporarily - which is in line with a bursty/stochastic SFH.
We compare the position in the SFR-stellar mass plot with more galaxies at similar redshifts (Fig. 9). The sSFR of E and W (51 and 38 Gyr-1 respectively) is similar to that of low-mass galaxies, but also to some very massive ones, such as REBELS-25 (sSFR=25 Gyr-1), a relaxed dynamically cold disc galaxy at similar redshift, â=â7.31 (Rowland et al. 2024). A recent large compilation of NIRSpec prism data and stacked spectra has shown that the average galaxy becomes progressively more bursty with increasing redshift as well as with increasing sSFR (Langeroodi & Hjorth 2024). Based on this, B14-65666 seems slightly less bursty (but consistent within error bars) than the average galaxy at similar redshift and sSFR. UV-faint galaxies at redshifts from 6 to 8 from the Cosmic Evolution Early Release Science (CEERS) JWST/NIRCam survey also show a bursty star formation history (Endsley et al. 2023). Galaxies from the ALMA-ALPINE [CâII] survey at redshifts 4 to 6 (Le FĂšvre et al. 2020; Faisst et al. 2020) show normal star formation following the main sequence of galaxies at these redshifts (Speagle et al. 2014). LIRGs at low redshift, with typical stellar masses in the range âMâ and SFR 30âMââyr-1 (Bellocchi et al. 2022) are placed above the low- main sequence plane (Elbaz et al. 2011) and indicating that at high redshifts LIRGs are less massive but more compact.
@@ 4.9 B14-65666: a merger of starburst galaxies in the early Universe
The picture that emerges from our analysis is that B14-65666 is a system of two galaxies in the process of interacting and merging. We have identified one of the components (E) as an extremely compact source, characterized by a stellar mass surface density close to that of nuclei of low-z galaxies. On the other hand, the second component of the system (galaxy W) shows a more elongated structure with four fainter clumps (Sugahara et al. 2025), reminiscent of star-forming regions along the tidal tails produced in gas-rich interacting galaxies. A significant velocity offset of 175âkmâs-1 between the two galaxies supports the merger scenario. Furthermore, the estimated stellar masses for each component are on the order of ââMâ, three to four orders of magnitude higher than typical stellar masses of star-forming clumps in galactic disks (ââââMâ, Adamo et al. 2024; Mowla et al. 2024), reinforcing the interpretation that we are observing two galaxy-scale systems. The stellar mass ratio between the components is â2:1, potentially consistent with a major merger. Additionally, we find differences in metallicity between the two components, with values derived with the SED-fitting of 0.24âZâ for E and 0.31âZâ for W, in agreement with line diagnostics from NIRSpec IFS (Jones et al. 2024). Collectively, the presence of tidal features, velocity offset, distinct stellar populations and metallicities, and the high stellar masses all provide compelling evidence that B14-65666 is an example of a merging system of two galaxies. In addition, galaxies E and W are experiencing a bursty phase with a SFR well above that of the main-sequence star-forming galaxies of the same mass at those redshifts.
@@
Recent simulations of the formation of galaxies during the Epoch of Reionization (Nakazato et al. 2024 and references therein) show that systems identified as clumpy, i.e. with two or more high surface density star-forming clumps as traced by emission lines, represent 10 of all simulated sources. Moreover, most of the clumpy systems are produced by mergers, with the young stellar populations forming along the gas-rich tidal tails, and with specific SFR of the order of 50 Gyr-1. However, these clumpy systems appear to be temporal and short-lived, merging into a single galaxy in about 50 Myr (Hashimoto et al. 2023). Under this scenario, the compact, and highly dense, galaxy E could already be a protobulge produced by the merger of previous clumps while galaxy W would be in the process of merging of smaller, less massive, clumps before the two main components, E and W, merge together in a single, more massive galaxy at a later time, i.e. lower redshifts (see Figure 5 in Nakazato et al. 2024).
@@
B14-65666 is not a unique system at those redshifts. Recent JWST observations have discovered merging systems of star-forming galaxies such as the interacting system SPT0311-58, which resides in the core of an extremely massive protocluster at â=â6.9 (Arribas et al. 2024), the source A2744-z7p9OD, the most distant protocluster at â=â7.88 (Hashimoto et al. 2023), or the JWSTâs Quintet, a merger of at least five galaxies containing over 17 clumps, at â=â6.7 (Hu et al. 2025). Detection and further studies with JWST of new systems like these will allow us to understand the formation process of massive galaxies in the early Universe.

### v2

@@ 4.5 Kinematics of the ionized ISM
The kinematics of the ionized gas can be measured with the H emission-line profile in our MRS data. Based on the integrated H line, we measure a redshift of â=â7.1513ââ0.0007 for the B14-65666 system, compatible with the values obtained with the FIR emission lines [OâIII]â88âm and [CâII]â158âm (â=â7.1521ââ0.0004, Hashimoto et al. 2019) and optical [OâIII] line (â=â7.1517ââ0.0001, Jones et al. 2024). The two galaxies of the B14-65666 system have a relative velocity of 175ââ28âkmâs-1 along our line of sight as derived from the centroid of the H line emission, showing it is a galaxy merger.
@@
The kinematics of the ionized gas traced by H appears to show differences between the two galaxies. Specifically, we find that the FWHM for galaxy W is higher than for galaxy E, with values of 312ââ44âkmâs-1 and 243ââ41âkmâs-1, respectively. These values agree within the uncertainties (from those measured) with the far-infrared emission lines, [CâII]â158âm and [OâIII]â88âm. The [CâII]â158âm emission has similar kinematics in both galaxies (FWHM of 288ââ41âkmâs-1 for galaxy W and 276ââ34âkmâs-1 555The FWHM value for clump E has been recalculated and differs from that reported in Hashimoto et al. (2019), as confirmed through a private communication with the authors. for galaxy E). For [OâIII]â88âm, galaxy E shows a broader line (322ââ41âkmâs-1) than galaxy W (267ââ42âkmâs-1). In Fig. 6, we present the comparison of these three lines for galaxies E+W and E and W. The observed differences in the two galaxies indicate the complex structure of the velocity field in the different phases of the ISM. While the H line traces the overall kinematics of the ionized gas, the [CâII]â158âm line is closely associated with the neutral gas and the [OâIII]â88 m with the high excitation gas. Simulations also show that different tracers are sensitive to different kinematic properties (Kohandel et al. 2024). In addition, the continuum imaging (Sugahara et al. 2025) of both galaxies shows a different morphology; galaxy E as a very compact source, while galaxy W shows a more extended structure, in particular in the rest-UV light, reminiscent of tidal tails. Moreover, low-spectral-resolution (Rââ1000) integral-field spectroscopy with NIRSpec suggests the presence of a narrow and a broad kinematic component in both galaxies and in all optical emission lines tracing the ionized gas (Jones et al. 2024). These two components are potentially associated with the motions in the host galaxy (narrow) and the presence of high-velocity gas outflows (broad). The width of the narrow component is only 144ââ11âkmâs-1 for galaxy E and 227ââ6âkmâs-1 for galaxy W. These are significantly narrower than the values derived from our H line and far-infrared emission lines, obtained with a better spectral resolution but lower S/N. We tested whether the observed H line profile could be compatible with the combination of the narrow and broad (653ââ10âkmâs-1 for galaxy E and 755ââ23âkmâs-1 for galaxy W) line components identified with NIRSpec. Our tests (see Fig. 7) indicate that the narrow+broad (N+B) component could be slightly compatible with our observed H line profiles. However, the simulated H N+B line profile appears to have a flux excess in the spectral range of the line (i.e., residuals mostly below zero) for galaxy E, while the faint broad component in galaxy W is close in velocity to the peak emission in galaxy E. Since the two galaxies are separated by only 0.44 arcsec, this secondary emission in W could be due to a contamination from the brighter emission from E (Jones et al. 2024). Further higher spectral resolution data with NIRSpec, and higher S/N MRS spectra, would be required to confirm the presence of such a broad line emission in the two galaxies.
@@
The consistency of the trends between our results and the ancillary data provides robust evidence of the dynamic differences between galaxies E and W. The broader profiles observed in galaxy W for both H and [CâII]â158âm may indicate enhanced turbulence or kinematic effects, potentially driven by interactions or feedback processes, as well as a possibility of a connection with outflows. On the other hand, the higher FWHM in [OâIII]â88 m for galaxy E suggests that highly ionized gas dynamics differ from those of lower ionization states, pointing to localized physical conditions or energetic phenomena. Also, dust attenuation could impact the interpretation, in particular in galaxy E where a large visual extinction is measured. The optical lines are more susceptible to dust extinction, which can affect measurements of line widths and kinematic features, whereas far-infrared lines are less impacted by such effects.
@@ 4.6 Stellar and dynamical mass
Although the line profiles of the different emission lines suggest that the gas is not fully (dynamically) relaxed, here we obtain an upper limit of the dynamical mass of each of the galaxies assuming the system is close to relaxation and characterized by the random motions traced by the ionized gas velocity dispersion; i.e., not dominated by rotational motions. Under this working hypothesis, we followed previous works on similar systems such as the low-z luminous infrared galaxies (e.g., Eq. (1) in Bellocchi et al. 2013). The dynamical masses we derive are (1.5ââ0.3)âââ and (8.5ââ2.4)âââ for galaxies E and W, respectively, assuming velocity dispersions of 103ââ17 kmâs-1 and 132ââ19âkmâs-1 (obtained with the FWHM from our H line), and the effective radii derived from our NIRCam images (see Sect. 3.4 for details). The total stellar masses are derived from the multiwavelength SED fitting and correspond to (15ââ9)âââMâ and (8ââ4)ââ108âMâ for galaxies E and W. Thus, while there is a good agreement between the stellar and dynamical masses for galaxy E, an order of magnitude difference appears for galaxy W. The difference in galaxy W could be due to the presence of a dominant molecular gas component in this galaxy amounting to close to 90% of the total mass. However, the stellar structure shows the presence of extended emission consisting of star-forming clumps along tidal tails (Sugahara et al. 2025). This suggests that this galaxy is far from relaxed, and therefore the estimated dynamical mass is highly uncertain.
@@ 4.7 Mass-size relation
The two galaxies of the B14-65666 system occupy a very different location in the stellar mass-size plane (see Fig. 8). Galaxy W, with an effective radius of 348 pc and a stellar mass of (8ââ4) 108 Mâ, is consistent with the expected values according to the relations derived for high-redshift galaxies from the JADES program (redshifts 4 to 10; Langeroodi & Hjorth 2023) and from the CEERS, PRIMER-UDS, and PRIMER-COSMOS surveys (redshifts 3 to 9; Allen et al. 2025). Therefore, galaxy W appears as an average galaxy at approximately redshift 7. On the other hand, the second component of the system, galaxy E, appears to be deviating from the standard relations, showing a very small and compact size (upper limit of 63 pc to the effective radius) for a stellar mass of 15 (â9)ââ108âMâ. This translates into a high-stellar-mass surface density (6âââMââpc-2 for E, versus 1â Mââpc-2 for W), close to the limit of the highest values (âMââpc-2) measured in the stellar clusters and nuclei of low-z galaxies (GrudiÄ et al. 2019). Galaxies as compact as galaxy E and with similar stellar surface densities are not uncommon at high- (Ălvarez-MĂĄrquez et al. 2025). Star-forming clumps and galaxies with an effective radius in the 30 to 200 pc range and a wide mass range (4ââ107 to 2ââ1010 Mâ) are already known at redshifts above 6 (see Fig. 8; Claeyssens et al. 2023; Baggen et al. 2023). Thus, B14-65666 is identified as a system that consists of a pair of interacting galaxies, with similar stellar masses (within a factor of 2) but with one of the galaxies (E) being extremely compact and already showing stellar-mass surface densities close to those measured in the nuclei of galaxies at low redshifts. The galaxies of the B14-65666 system are much less massive (ââ20âââ30) and far more compact (ââ10âââ60) than the main-sequence and most luminous star-forming galaxies at low redshifts, with effective radii â3.5 kpc (Leroy et al. 2021) and â2.2 kpc (Bellocchi et al. 2022), respectively (see Fig. 8). Thus, galaxies E and W appear to be building up their stellar mass as they go through the advanced phases of the interaction/merging process.
@@ 4.8 Starburst nature
The analysis of the SFR derived with the H-versus-stellar-mass plane reveals that both galaxies E and W are well above the main sequence (MS) of galaxies at redshift 7 (Heintz et al. 2023). Galaxies E and W of our system are situated between the region defined by Caputi et al. (2017) for H emitters, the starburst (SB) region, and the MS (see Fig. 9). Such extremely compact starburst galaxies are expected to have short depletion times (ÂĄ100 Myr, see Franco et al. 2020), and therefore galaxy E may quickly quench out of gas exhaustion âat least temporarilyâ which is in line with a bursty/stochastic SFH.
We compared the position in the SFR stellar-mass plot with more galaxies at similar redshifts (Fig. 9). The sSFRs of E and W (51 and 38 Gyr-1, respectively) are similar to those of low-mass galaxies, but also to some very massive ones, such as REBELS-25 (sSFR=25 Gyr-1), a relaxed, dynamically cold disc galaxy at a similar redshift, â=â7.31 (Rowland et al. 2024). A recent large compilation of NIRSpec prism data and stacked spectra showed that the average galaxy becomes progressively more bursty with increasing redshift, as well as with increasing sSFR (Langeroodi & Hjorth 2024). Based on this, B14-65666 seems slightly less bursty (but consistent within error bars) than the average galaxy at similar redshift and sSFR. UV-faint galaxies at redshifts from 6 to 8 from the Cosmic Evolution Early Release Science (CEERS) JWST/NIRCam survey also show a bursty star formation history (Endsley et al. 2023). Galaxies from the ALMA-ALPINE [CâII] survey at redshifts 4 to 6 (Le FĂšvre et al. 2020; Faisst et al. 2020) show normal star formation following the MS of galaxies at these redshifts (Speagle et al. 2014). LIRGs at low redshifts, with typical stellar masses in the âMâ and SFR 30âMââyr-1 range (Bellocchi et al. 2022), are placed above the low- MS plane (Elbaz et al. 2011) and indicate that at high redshifts LIRGs are less massive but more compact.
@@ 4.9 B14-65666: A merger of starburst galaxies in the early Universe
The picture that emerges from our analysis is that B14-65666 is a system of two galaxies in the process of interacting and merging. We identified one of the components (E) as an extremely compact source, characterized by a stellar-mass surface density close to that of the nuclei of low-z galaxies. On the other hand, the second component of the system (galaxy W) shows a more elongated structure with four fainter clumps (Sugahara et al. 2025), reminiscent of star-forming regions along the tidal tails produced in gas-rich interacting galaxies. A significant velocity offset of 175âkmâs-1 between the two galaxies supports the merger scenario. Furthermore, the estimated stellar masses for each component are on the order of ââMâ, which is three-to-four orders of magnitude higher than typical stellar masses of star-forming clumps in galactic disks (ââââMâ, Adamo et al. 2024; Mowla et al. 2024), reinforcing the interpretation that we are observing two galaxy-scale systems. The stellar-mass ratio between the components is â2:1, which is potentially consistent with a major merger. Additionally, we find differences in metallicity between the two components, with values derived using the SED-fitting of 0.24âZâ for E and 0.31âZâ for W; this is in agreement with line diagnostics from NIRSpec IFS (Jones et al. 2024). Collectively, the presence of tidal features, velocity offset, distinct stellar populations and metallicities, and the high stellar masses all provide compelling evidence that B14-65666 is an example of a merging system of two galaxies. In addition, galaxies E and W are experiencing a bursty phase with an SFR well above that of the main-sequence star-forming galaxies of the same mass at those redshifts.
@@
Recent simulations of the formation of galaxies during the EoR (Nakazato et al. 2024 and references therein) show that systems identified as clumpy âi.e., with two or more high-surface-density star-forming clumps as traced by emission linesâ represent 10 of all simulated sources. Moreover, most of the clumpy systems are produced by mergers, with the young stellar populations forming along the gas-rich tidal tails, and with specific SFR of the order of 50 Gyr-1. However, these clumpy systems appear to be temporal and short-lived, merging into a single galaxy in about 50 Myr (Hashimoto et al. 2023). Under this scenario, the compact, and highly dense, galaxy E could already be a protobulge produced by the merger of previous clumps, while galaxy W would be in the process of merging of smaller and less massive clumps before the two main components, E and W, merge together in a single, more massive galaxy at a later time; i.e., at lower redshifts (see Figure 5 in Nakazato et al. 2024).
@@
B14-65666 is not a unique system at those redshifts. Recent JWST observations show merging systems of star-forming galaxies such as the interacting system SPT0311-58, which resides in the core of an extremely massive protocluster at â=â6.9 (Arribas et al. 2024); the source A2744-z7p9OD, the most distant protocluster at â=â7.88 (Hashimoto et al. 2023); or the JWSTâs Quintet, a merger of at least five galaxies containing over 17 clumps, at â=â6.7 (Hu et al. 2025). The detection with the JWST and further studies of new systems such as these will allow us to understand the formation process of massive galaxies in the early Universe.